Body A weighs 108 N, and body B weighs 35 N. The coefficients of friction between...

uploaded imageBody A weighs 108 N, and body B weighs 35 N. The coefficients of friction between A and the incline are μs = 0.56 and μk = 0.25. Angle θ is 40°. Let the positive direction of an x axis be up the incline.
(a) Find the acceleration of A if A is initially at rest.
( ____ m/s2) ihat

(b) Find the acceleration of A if A is moving up the incline.
( ____ m/s2) ihat

(c) Find the acceleration of A if A is moving down the incline.
( ____ m/s2) ihat
Answer #1
weight of the body A , WA = mAg = 108 N
mass of the block A, mA = (108 N)/(9.8 m/s2)
                                      = 11.02 kg
weight of the body B , WB =mB g = 35 N
mass of the block B, mA = (35 N)/(9.8 m/s2)
                                      = 3.57 kg
The coefficients of friction between A and the incline are
         μs = 0.56 and μk = 0.25.
angle θ = 400
.......................................................................
a)
if the body A is initially at rest :
using Newton's second law of motion ,
T - f - mAgsinθ = 0
and   mBg - T = 0
        T = mBg
hence ,
mBg - f - mAgsinθ = 0
frictional force f = mBg - (mAg)sinθ
                         = (35 N) - (108 N)sin40
                         = -34.42 N
normal force N = (mAg) cosθ
                        = (108 N)coas40
                        = 82.73 N
so , maximum frictional force fs,max   = μsN
                                                        = (0.56)(82.73 N)
                                                        = 46.33 N
from this ,
          f < fs,max
so , the acceleration a = 0 m/s2 i^
..........................................................................
b)
if A is moving up the incline :
using Newton's second law of motion ,
           mAa = T - fk - mAgsinθ    ............... (1)
    here , frictional force fk = μkN
                                        = μkmAg cosθ
equation (1) becomes
         mAa = T - μkmAg cosθ - mAgsinθ     ............. (2)  
for block B :
       mBa = mBg - T
tension T = mBg - mBa   ................ (3)
substitute the eq (3) in eq (2) , we get
       mAa = mBg - mBa - μkmAg cosθ - mAgsinθ
therefore , acceleration is
        a = [(mBg) - (mAg) sinθ - μk(mAg) cosθ] /(mA + mB)      ............. (4)   
substitute the given data in eq (4) , we get
       a = [(35 N) - (108 N) (sin40) - (0.25)(108 N) (cos40)] /( 11.02 kg+ 3.57 kg)
          = (-3.77 m/s2) i^ [objects are slowing down]
...............................................................
c)
if A is moving down the incline :
using Newton's second law of motion ,
           mAa = T + fk - mAgsinθ    ............... (5)
    here , frictional force fk = μkN
                                        = μkmAg cosθ
equation (5) becomes
         mAa = T + μkmAg cosθ - mAgsinθ     ............. (6)  
for block B :
       mBa = mBg - T
tension T = mBg - mBa   ................ (7)
substitute the eq (7) in eq (6) , we get
       mAa = mBg - mBa  + μkmAg cosθ - mAgsinθ
therefore , acceleration is
        a = [(mBg) - (mAg) sinθ + μk(mAg) cosθ] /(mA + mB)      ............. (8)   
substitute the given data in eq (4) , we get
       a = [(35 N) - (108 N) (sin40) + (0.25)(108 N) (cos40)] /( 11.02 kg+ 3.57 kg)
          = (-0.94 m/s2) i^ [objects are speeding up]